OpenStudy (anonymous):

Can someone give me an equation that is two steps. It needs to have a solution of all real numbers

Directrix (directrix):

x + 2 = x + 2 x = x

OpenStudy (anonymous):

it needs to be two step @Directrix

Directrix (directrix):

I counted the problem as the first step. --------------------------- How about this: 3x + 2 = 3x + 2 3x = 3x x = x @MiaS